Q1 Question 1 1 Point d(x, y) = (x – y| is a metric on R. O true O false Save Answer Q2 Question 2 1 Point The Euclidean distThank you!
Answer #1

Let A be arnon-empty set. is said to be matrie on A. A funesion di AXANIR it Satisly following condenon- ņ daly) 7, O , AnneFuelidean distance fromuler, d (y) = key-mg2+ (n = 132 - + (-3)? where (*1,*23*--, an), y=(4,,2,, ym) from the definasion ofxongitor 2) izn 2 الا ainy :) Y:) + 2 (93-9;) (03-21) +(2,0-2,0) 2 = ? (q; ;) 2212 2:07:) (4:-) + 2 (y; -21) 124 al (1) Is03 Come norm on the veter Brezy Matrie on a vector space is deferrmine by word, d is a matre on a vector space for some norm